A and B are independent events. Use the following probabilities to answer the question. Round to 4 decimal places.
P(A) = 0.75, P(A and B) = 0.3, find P(B)


Answer by ikleyn(52848)

.

Since A and B are independent,


    P(A and B) = P(A)*P(B),

or

    0.3 = 0.75*P(B).


It gives  P(B) = 0.3/0.75 = 0.4.    ANSWER

Solved.